# KnowEEG: Explainable Knowledge Driven EEG Classification

### v1

@@ 4.6. Benefits and Limitations
The main benefit of KnowEEG is the explainability of the model which is in contrast to many state of the art deep learning models. The features themselves have meaning in the generalized time series or EEG domain and the Fusion Forest is a tree-based model that allows users to directly access feature importances. We demonstrate this benefit in 4.5. We show that for the Crowdsourced dataset our explainable pipeline can allow users to discover correct knowledge about the two classes. Notably, we find Occipital Lobe signal statistics and alpha coherence to be discriminative of the eyes open and eyes closed classes. Both findings are confirmed by existing neuroscience literature. Another benefit of our pipeline is high performance across a diverse set of EEG classification tasks. Versus 7 state-of-the-art baselines on five datasets on 10 performance metrics our pipeline is best on 6/10 metrics, second on 3/10 and outside of the top two models on only 1/10 metrics. Overall, KnowEEG is the best performing model across 5 different EEG datasets. Finally, our pipeline has the significant benefit of not requiring GPU resources for training. The most computationally expensive part of our pipeline is calculation of features that is done on CPU. This means KnowEEG requires fewer specialized resources versus competitor state-of-the-art deep learning models and is therefore more accessible.
@@
4.7. Future Work
Future work will explore additional EEG datasets to determine if the performance of KnowEEG remains high. This should involve Brain Computer Interface (BCI) EEG datasets such as those discussed in (bci, ). BCI datasets broadly address the task of enabling a person’s brain signals to communicate with external devices which is distinctly different from the five datasets already explored.
KnowEEG should be deployed on real world data in the healthcare domain to determine if the pipeline can successfully discriminate between classes where the differences between the classes are not already understood. Neurodegenerative diseases such as Alzheimers and Parkinsons’ would be a good use case. KnowEEG could potentially uncover previously unknown difference in EEG activity between healthy and diseased individuals furthering advancements in this field.
Further analysis in future work should be done on why KnowEEG outperforms competitors. For example, one could determine what the extent of the shared information between learnt self-supervised representations such as in EEG2Rep (mohammadi2024eeg2rep, ) and the features in KnowEEG. This could aid in the development of higher performance models in the future.
5. Conclusion
We present KnowEEG, an explainable high-performance pipeline for EEG classification. We demonstrate through extensive experimentation across five different EEG classification tasks that KnowEEG outperforms state-of-the-art deep learning methods. KnowEEG achieves an average rank of first in two performance metrics (accuracy / balanced accuracy and AUROC / W-F1) versus competitor models. KnowEEG has the added benefits of not requiring GPUs for training and is also explainable.
We illustrate the explainability properties of KnowEEG in Section 4.5, using the Crowdsource dataset as an example. With directly accessible feature importances and interpretable features, KnowEEG allowed us to correctly discover knowledge about the two classes, specifically that the eyes closed state has increased alpha coherence versus eyes open and that statistics from the Occipital region electrodes were most discriminative of the two classes.
@@ Appendix A KnowEEG Set Up
The TSFresh (christ2018time, ) python package with version 0.20.3 was used to calculate the per electrode statistics with ’Efficient’ setting for feature calculation. The MNE python package (MNE_2013, ) version 1.7.1 was used to calculate the connectivity metrics, further details on connectivity metrics are provided in appendix B. The python package Scikit-learn (pedregosa2011scikit, ) with version 1.6.1 was used to generate the trees of the Fusion Forest. Decision trees with default hyperparameter settings were used for the Fusion Forest (criterion =’gini’, max depth=None, min samples split =2, min samples leaf = 1, max features = None).
Appendix B Connectivity Metrics
Connectivity metrics for KnowEEG were selected using literature and are not exhaustive (MNE_2013, ) (kuang2022phase, ) . The selected metric is a hyperparameter of the pipeline. Therefore users can add metrics to their pipeline that they expect to be informative.
@@ Appendix B Connectivity Metrics
The MNE package was used for all connectivity metric calculations (except for correlation) and can therefore be referred to for exact definitions (MNE_2013, ). We use the same names below as in the MNE package.
@@
Table 4. Overview of segmentation for signals for each dataset. Required in order to calculate specific connectivity metrics (Pairwise Phase Consistency, Phase Lag Value, Phase Lag Index, Directed Phase Lag Index, Weighted Phase Lag Index).
Appendix C Datasets
Datsets and preprocessing follow the same protocol as in (mohammadi2024eeg2rep, ).
C.1. Emotiv datasets
@@
The DREAMER (katsigiannis2017dreamer, ) dataset is a multimodal database containing both electroencephalogram (EEG) and electrocardiogram (ECG) signals. These signals are recorded during affect elicitation with audi-visual stimuli (katsigiannis2017dreamer, ). Data is recorded from 23 participants along with self-assessment of affective state after each stimuli. Self assessment is in terms of valence, arousal and dominance. For classification we use the arousal labels as per (mohammadi2024eeg2rep, ). We utilize the toolkit Torcheeg for preprocessing which consists of low-pass and high-filters. We do not use the ECG data and use only the EEG data. The DREAMER dataset can be accessed here [^1] https://zenodo.org/records/546113.
@@
The Crowdsourced (gomezram_eyes_op_close, ) dataset was recorded with participants at rest in the eyes closed or eyes open state. Each recording is 2 minutes in total. There are 60 total participants with only 13 recording data for both states. Data is recorded using 14-channel EPOC devices and is initially recorded at 2048Hz and then downsampled to 128Hz. This data can be accessed via the Open Science Framework  [^2] https://osf.io/9bvgh
@@
The STEW dataset (lim2018stew, ) is an open access EEG dataset for multitasking and mental workload analysis. There are 48 total subjects and data is recorded using a 14-channel Emotiv EPOC headset. Data is recorded for subjects at baseline (rest) and under workload in the SIMKAP mutitasking setting. EEG recordings are of length 2.5 minutes and recorded at 128Hz. Participants recorded their perceived mental workload on a scale of 1 to 9. STEW also has binary labels recording workload of above 4 as high and below 4 as low. These binary labels are used for our classification task. STEW is accessiible via the IEEE DataPort 3336https://ieee-dataport.org/open-access/stew-simultaneous-task-eeg-workloaddataset.
C.2. Temple University Hospital (TUH) Datasets
The TUH EEG Events dataset (TUEV) (TUEV_dataset, ) and the TUH Abnormal EEG corpus (TUAB) (TUAB_dataset, ) are accessible on request from the Neural Engineering Data Consortium (NEDC) here  [^4] https://isip.piconepress.com/projects/nedc/html/tuh_eeg/. The TUH datasets were processed in accordance with standard 16 EEG montages (mohammadi2024eeg2rep, ) and the 10-20 international system with: ”FP1-F7”, ”F7-T7”, ”T7-P7”, ”P7-O1”, ”FP2-F8”, ”F8-T8”, ”T8-P8”, ”P8-O2”, ”FP1-F3”, ”F3-C3”, ”C3-P3”, ”P3-O1”, ”FP2-F4”, ”F4-C4”, ”C4-P4”, and ”P4-O2”.
@@ C.2. Temple University Hospital (TUH) Datasets
The TUH EEG Events Corpus (TUEV) (TUEV_dataset, ) (mohammadi2024eeg2rep, ) contains EEG data with samples segmented into six categories. The categories are spike and sharp wave, eye movement, artifact, background generalized periodic epileptiform discharges and periodic lateralized epileptiform discharges (TUEV_dataset, ).
@@
The TUH Abnormal EEG Corpus (TUAB) (TUAB_dataset, ) (mohammadi2024eeg2rep, ) is a large-scale corpus of EEGs recordings designed to support research on automated EEG interpretation. This corpus is drawn from a diverse group of subjects with EEG labeled as either normal or abnormal (TUAB_dataset, ).
Appendix D Theta Coherence Plot
Figure 7 shows the mean theta Coherence for eyes closed vs. eyes open participants (training and validation sets). Though differences are present, with eyes closed again showing higher coherence - differences are visually less significant than the alpha band coherence plot (Figure 4) as implied by the feature importance scores.
[…]
Heatmap of the head showing theta coherence for eyes closed and eyes open. Refer to caption
Figure 7. Theta coherence across participants in training and validation set for eyes closed and eyes open plotted as heatmap on the head.

### v2

@@ 4.6 Benefits and Limitations
The main benefit of KnowEEG is the explainability of the model which is in contrast to many state of the art deep learning models. The features themselves have meaning in the generalized time series or EEG domain and the Fusion Forest is a tree-based model that allows users to directly access feature importances. We demonstrate this benefit in 4.5. We show that for the Crowdsourced dataset our explainable pipeline can allow users to discover correct knowledge about the two classes. Notably, we find Occipital Lobe signal statistics aswell as gamma band activity, alpha band activity and alpha band connectivity to be discriminative of the eyes open and eyes closed classes. Both findings are confirmed by existing neuroscience literature. Another benefit of our pipeline is high performance across a diverse set of EEG classification tasks. Versus 7 state-of-the-art baselines on five datasets on 10 performance metrics our pipeline is best on 5/10 metrics, second on 4/10 and outside of the top two models on only 1/10 metrics. Overall, KnowEEG is the best performing model across 5 different EEG datasets. Finally, our pipeline has the significant benefit of not requiring GPU resources for training. The most computationally expensive part of our pipeline is calculation of features that is done on CPU. This means KnowEEG requires fewer specialized resources versus competitor state-of-the-art deep learning models and is therefore more accessible.
@@
4.7 Future Work
Future work will explore additional EEG datasets to determine if the performance of KnowEEG remains high. This should involve Brain Computer Interface (BCI) EEG datasets such as those discussed in [11]. BCI datasets broadly address the task of enabling a person’s brain signals to communicate with external devices which is distinctly different from the five datasets already explored.
KnowEEG should be deployed on real world data in the healthcare domain to determine if the pipeline can successfully discriminate between classes where the differences between the classes are not already understood. Neurodegenerative diseases such as Alzheimers and Parkinson’s would be a good use case. KnowEEG could potentially uncover previously unknown difference in EEG activity between healthy and diseased individuals furthering advancements in this field.
Further analysis in future work should be done on why KnowEEG outperforms competitors. For example, one could determine what the extent of the shared information between learnt self-supervised representations such as in EEG2Rep [32] and the features in KnowEEG. This could aid in the development of higher performance models in the future.
5 Conclusion
We present KnowEEG, an explainable high-performance pipeline for EEG classification. We demonstrate through extensive experimentation across five different EEG classification tasks that KnowEEG outperforms state-of-the-art deep learning methods. KnowEEG achieves top or near-top performance across evaluation metrics (accuracy / balanced accuracy and AUROC / W-F1) versus competitor models. KnowEEG has the added benefits of not requiring GPUs for training and is also explainable.
Section 4.5 illustrates the explainability properties of KnowEEG using the Crowdsource dataset. By providing transparent feature importances and interpretable feature representations, KnowEEG allowed us to correctly identify established differences between the two states; reduced alpha synchronisation, elevated alpha-band power, and lower gamma activity during eyes-closed versus eyes-open with statistics from the Occipital-region electrodes most discriminative for classification.
@@ 5 Conclusion
6 Funding
This work was supported by UK Research and Innovation grant EP/S022937/1: Interactive Artificial Intelligence.
@@ Appendix A KnowEEG Set Up
The TSFresh [12] python package with version 0.20.3 was used to calculate the per electrode statistics with ‘Efficient’ setting for feature calculation. The MNE python package [18] version 1.7.1 was used to calculate the connectivity metrics, further details on connectivity metrics are provided in appendix C. The python package Scikit-learn [34] with version 1.6.1 was used to generate the trees of the Fusion Forest. Decision trees with default hyperparameter settings were used for the Fusion Forest (criterion =‘gini’, max depth=None, min samples split =2, min samples leaf = 1, max features = None).
Appendix B Functional Power Connectivity
In addition to correlation and conventional phase / coherence based connectivity measures we introduce an amplitude-based functional connectivity set of features which we term Functional Power Connectivity (FPC). For each electrode and frequency band (delta, theta, alpha, sigma, beta, gamma) we compute relative bandpower. We then define FPC features as the pairwise log-ratios of relative bandpower between electrodes.
Therefore for each band between each electrode pair i,j we compute :
$\mathrm{FPC}_{i,j}=\log P^{\mathrm{rel}}_{i}-\log P^{\mathrm{rel}}_{j}=\log\left(\frac{P^{\mathrm{rel}}_{i}}{P^{\mathrm{rel}}_{j}}\right)$ (1)
This means for a set of 14 electrodes there are in total 630 total FPC features comprising 6 x 14 band power features plus 91 x 6 bands of electrode-electrode features. These features capture regional band power as well as spatial differences in oscillatory power across regions which can be interpreted as power-based functional connectivity markers.
Appendix C Connectivity Metrics Overview
Connectivity metrics for KnowEEG were selected using literature and are not exhaustive [18] [26] . The selected metric is a hyperparameter of the pipeline. Therefore users can add metrics to their pipeline that they expect to be informative.
@@ Appendix C Connectivity Metrics Overview
The MNE package was used for all connectivity metric calculations (except for correlation) and can therefore be referred to for exact definitions [18]. We use the same names below as in the MNE package.
@@
  • 2.
    Functional Power Connectivity
Selected Metrics (with signal segmentation) are:
  • 1.
@@
Table 4: Overview of segmentation for signals for each dataset. Required in order to calculate specific connectivity metrics (Coherence, Imaginary Coherence, Pairwise Phase Consistency, Phase Lag Value, Phase Lag Index, Directed Phase Lag Index, Weighted Phase Lag Index).
Appendix D Datasets
Datasets and preprocessing follow the same protocol as in [32].
D.1 Emotiv datasets
@@
The DREAMER [24] dataset is a multimodal database containing both electroencephalogram (EEG) and electrocardiogram (ECG) signals. These signals are recorded during affect elicitation with audi-visual stimuli [24]. Data is recorded from 23 participants along with self-assessment of affective state after each stimuli. Self assessment is in terms of valence, arousal and dominance. For classification we use the arousal labels as per [32]. We utilize the toolkit Torcheeg for preprocessing which consists of low-pass and high-filters. We do not use the ECG data and use only the EEG data. The DREAMER dataset can be accessed here [^1] https://zenodo.org/records/546113.
@@
The Crowdsourced [17] dataset was recorded with participants at rest in the eyes closed or eyes open state. Each recording is 2 minutes in total. There are 60 total participants with only 13 recording data for both states. Data is recorded using 14-channel EPOC devices and is initially recorded at 2048Hz and then downsampled to 128Hz. This data can be accessed via the Open Science Framework  [^2] https://osf.io/9bvgh
@@
The STEW dataset [28] is an open access EEG dataset for multitasking and mental workload analysis. There are 48 total subjects and data is recorded using a 14-channel Emotiv EPOC headset. Data is recorded for subjects at baseline (rest) and under workload in the SIMKAP mutitasking setting. EEG recordings are of length 2.5 minutes and recorded at 128Hz. Participants recorded their perceived mental workload on a scale of 1 to 9. STEW also has binary labels recording workload of above 4 as high and below 4 as low. These binary labels are used for our classification task. STEW is accessiible via the IEEE DataPort 3336https://ieee-dataport.org/open-access/stew-simultaneous-task-eeg-workloaddataset.
D.2 Temple University Hospital (TUH) Datasets
The TUH EEG Events dataset (TUEV) [21] and the TUH Abnormal EEG corpus (TUAB) [29] are accessible on request from the Neural Engineering Data Consortium (NEDC) here  [^4] https://isip.piconepress.com/projects/nedc/html/tuh_eeg/. The TUH datasets were processed in accordance with standard 16 EEG montages [32] and the 10-20 international system with: ”FP1-F7”, ”F7-T7”, ”T7-P7”, ”P7-O1”, ”FP2-F8”, ”F8-T8”, ”T8-P8”, ”P8-O2”, ”FP1-F3”, ”F3-C3”, ”C3-P3”, ”P3-O1”, ”FP2-F4”, ”F4-C4”, ”C4-P4”, and ”P4-O2”.
@@ D.2 Temple University Hospital (TUH) Datasets
The TUH EEG Events Corpus (TUEV) [21] [32] contains EEG data with samples segmented into six categories. The categories are spike and sharp wave, eye movement, artifact, background generalized periodic epileptiform discharges and periodic lateralized epileptiform discharges [21].
@@
The TUH Abnormal EEG Corpus (TUAB) [29] [32] is a large-scale corpus of EEGs recordings designed to support research on automated EEG interpretation. This corpus is drawn from a diverse group of subjects with EEG labeled as either normal or abnormal [29].